OpenStudy (anonymous):

What is the empirical formula of a compound that is 27.3% C and 72.7% O?

OpenStudy (anonymous):

@vanessanne

OpenStudy (nincompoop):

this should be easy stoichiometry :)

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

that it what u were trying to said hopefully something instead leave that comment

OpenStudy (anonymous):

Assume that there is 100 gr of the compound so we got 27.3 gr C and 72.7 gr O \[?mol C = 27.3 gr C \times \frac{ 1 mol C }{12 gr C } = 2.275 mol C\] \[?mol O = 72.2 gr O \times \frac{ 1mol O }{16 gr O } = 4.5125 mol O\] \[4.5125 \div 2.275 \approx 2\] so the empirical formula would be CO2